OpenStudy (anonymous):

Determine the value of dy/dt at x = 2 when y = 3x^2− 5x and dx/dt = 3.

OpenStudy (campbell_st):

you need to differentiate.. but it looks like it needs to be dy/dx and not dy/dt the general rule is if \[y = x^n.... then... \frac{dy}{dx} = n \times x^{n -1}\] using that rule you will have \[\frac{dy}{dx} = 3 \times 2 x^{2 -1} - 5 \times 1 \times x^{1-1}\] this simplifies to \[\frac{dy}{dx} = 6x - 5\] so for the 1st part substitute x = 2 into the above equation and solve for dy/dx The next part is let dy/dx = 3 and solve for x so you will have 3 = 6x - 5 solve for x hope it helps